OpenStudy (anonymous):

What is the volume of the solid of revolution generated by revolving the area bounded by y = 2, y = x+3 and x = 4 around the x-axis

OpenStudy (amistre64):

ahhh volume of solids :) nice

OpenStudy (amistre64):

isnt that just a cone?

OpenStudy (anonymous):

Lol - up for helping me out one more time?

OpenStudy (amistre64):

spin this right?

OpenStudy (amistre64):

it amounts to a cylinder and a cone :)

OpenStudy (anonymous):

Yes, that's the same sketch I made

OpenStudy (anonymous):

Lol

OpenStudy (amistre64):

y = 2 when x = -1; -1 to 4 is 5 units; area of the left part is 5(pi) 2^2

OpenStudy (amistre64):

20 pi is the area of the "cylindar"; now add the cone...

OpenStudy (amistre64):

we have a base area of pi 2^2 = 4pi; and a height from -1 to -3 = 2

OpenStudy (amistre64):

4pi(2) ----- = 8pi/3 add the other 3 8pi + 60pi --------- = 68pi/3 3

OpenStudy (amistre64):

thats my guess :) cone plus cylinder = total;

myininaya (myininaya):

this might be wrong but I have to leave

OpenStudy (amistre64):

that might be right :)

OpenStudy (anonymous):

hmm -

OpenStudy (amistre64):

we can solve both ways ;)

myininaya (myininaya):

you guys have fun i have a date lol

OpenStudy (anonymous):

hmm - i'm not sure if 68pi/3 is right

OpenStudy (amistre64):

we move everything to the right by 1 to get bounbes from 0 to 5; then rotate the triangle in a torus

OpenStudy (anonymous):

It was a multiple choice question from the book - but no answer in the back to confirm

OpenStudy (amistre64):

\[\pi \int\limits_{0}^{5} [f(x)]^2 dx\]

OpenStudy (amistre64):

y=x+3 when moved to the right by 1 becomes: y+1 = x+3 y = x+2 integrate (x+2)^2

OpenStudy (amistre64):

(x+2)^3 ------- = F(x); play with the 5 and we get... 343pi/3 3

OpenStudy (anonymous):

the numerator seems to be off - it's either 14 pi, 275 pi or 140 pi

OpenStudy (anonymous):

I can't seem to get any of those numbers on accident either!

OpenStudy (amistre64):

lets see if I integrated it forrectly :) u = x+2; du = dx [S] u^2 du -> (u^3)/3 = (x+2)^3/3 thats right....

OpenStudy (amistre64):

its from 0 to 5 so..... thats where I misssed it lol pi(343-8)/3

OpenStudy (amistre64):

i assumed 0 was just 0

OpenStudy (amistre64):

we have to subtract something; the area of the cylindar beneath it.....

OpenStudy (amistre64):

we made a solid and forgot to ...by we I mean me lol..... forgot to take away the solid formed beneath it

OpenStudy (anonymous):

lol

OpenStudy (amistre64):

subtract another 20pi

OpenStudy (amistre64):

pi(343-68)/3 275pi/3

OpenStudy (anonymous):

Great! Glad we (I mean you) could figure it out and explain it to me! Lol

OpenStudy (amistre64):

lol.... i took the whole thing and forgot to cut out the center of the doughnot.... that was my error the first go araound

OpenStudy (anonymous):

No worries!! I'm glad you could figure it out!

OpenStudy (anonymous):

I appreciate the help very much! I may post a few more if you're still around!

OpenStudy (amistre64):

this is the visual lol

OpenStudy (amistre64):

i gotsta take a macroeconomics final so good luck ;)

OpenStudy (anonymous):

Ahh! Haha

OpenStudy (anonymous):

Good luck!